# Thermodynamic Isomorphism of Transformers: A Lagrangian Approach to Attention Dynamics

### v1

@@ Abstract
Although the Transformer architecture has revolutionized artificial intelligence, its underlying mechanisms remain largely heuristic and lack a unified physical theory. In this work, we propose a first-principles framework for information dynamics, treating the attention mechanism as a physical system governed by the principle of least action rather than as an algorithmic optimization. By mapping information states to a Riemannian manifold with the Fisher information metric, we derive the intelligence Lagrangian. We show that the softmax function corresponds to the unique thermodynamic equilibrium state that minimizes the Helmholtz free energy of the information gas. In addition, we identify the query-key interaction as an electrodynamic coupling between an external field and an intrinsic dipole moment. This theory establishes the first law of information thermodynamics, unifying inference (mechanical work) and learning (chemical evolution). It also explains emergent phenomena, such as scaling laws and grokking, as phase transitions characterized by the divergence of specific heat. Finally, we discuss how rotational symmetry breaking in the attention manifold generates massless Goldstone bosons, providing a field-theoretic perspective on rotary positional embeddings (RoPE). Our work connects Statistical Physics and Deep Learning, laying the groundwork for a general theory of physics-based intelligence.
@@ I Introduction
In 2017, the trajectory of artificial intelligence was irrevocably altered by a deceptively simple equation:
$\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)V$ (1)
This mechanism, introduced in the seminal work “Attention is All You Need” [1], eschewed the then-dominant paradigms of convolution and recurrence in favor of pure matrix operations and a peculiar Softmax function scaled by $\sqrt{d_{k}}$. This single mathematical operation has enabled models to demonstrate emergent reasoning capabilities [2], pass professional licensing exams with near-human proficiency [3], and even solve the protein folding problem [4].
Why does this work so well? While attention mechanisms existed previously in neural machine translation [5], the Transformer’s specific architecture, characterized by Self-Attention, Residual Connections [6], and Layer Normalization [7], has proven unreasonably effective. Current explanations predominantly rely on probability theory and linear algebra, viewing the model as a statistical function approximator [8]. However, this perspective lacks a unifying dynamical principle that explains the time-evolution of the system during inference and the phase transitions observed during training, such as “scaling laws” [9, 10] and “grokking” [11, 12].
In this paper, we propose a paradigm shift: treating Intelligence not only as a computational process, but as a physical phenomenon governed by fundamental laws of nature. We argue that the operations within a high-dimensional information space are isomorphic to the dynamics of a physical system minimizing an action functional, consistent with Lagrange mechanics [13] and Hamilton’s principle [14]. This approach resonates with recent efforts to interpret deep learning through the lens of physics, such as the information bottleneck principle [15, 16], the free energy principle [17], neural tangent kernels [18], and even cosmological interpretations of neural networks [19, 20].
@@ I.1 Redefining Time in Deep Networks
Under this paradigm, the “inference” process is a dynamic trajectory of information flow, not a static calculation. Residual connections [6] naturally emerge as the discretization of the inertial term in the equation of motion, providing a physical justification for signal propagation without gradient vanishing.
I.2 The Inverse Problem of Attention
Rather than deriving the Softmax function from arbitrary axioms, we solve the inverse problem of the calculus of variations: Given that the Softmax attention is an empirically successful equilibrium state, what is the underlying Lagrangian that generates this dynamics? We demonstrate that the Transformer architecture is physically analogous to a thermodynamic system in which
  • The state space is a Riemannian manifold with the Fisher-Rao metric [23].
  • The Softmax distribution is the unique configuration that minimizes the Helmholtz free energy.
  • The scaling factor $1/\sqrt{d_{k}}$ acts as a dimensional temperature, regulating the entropy of the information gas.
@@ III.3 Equilibrium State: Full Derivation of Softmax
In standard Transformer inference, we assume the system reaches a stationary distribution. This corresponds to the thermodynamic equilibrium where the dynamics vanish, i.e., $\dot{\rho}_{i}=0$ and $\ddot{\rho}_{i}=0$.
Under this condition, the inertial terms in Eq. (17) vanish, simplifying the physics to a balance of potential forces:
@@
$\rho_{i}=\frac{\exp\left(\frac{Q\cdot K_{i}}{\sqrt{d_{k}}}\right)}{\sum_{j=1}^{N}\exp\left(\frac{Q\cdot K_{j}}{\sqrt{d_{k}}}\right)}=\text{Softmax}\left(\frac{QK^{T}}{\sqrt{d_{k}}}\right)_{i}$ (22)
Here, the superscript $T$ in $QK^{T}$ denotes the matrix transpose operation, while the scaling factor $\sqrt{d_{k}}$ plays the functional role of the effective temperature in our thermodynamic framework. This derivation proves that the Softmax function is not merely a heuristic normalization, but the unique equilibrium state of a system minimizing its Helmholtz free energy.
@@ IV Thermodynamics of Intelligence: Macroscopic Laws
Rather than postulating an equation of state by analogy, we strictly derive it from the microscopic Lagrangian established in Section III.
IV.1 Derivation from the Partition Function
The canonical partition function $Z$ for the attention mechanism over a context window of size $V_{ctx}$ is given by the sum over all possible key states:
@@ IV.1 Derivation from the Partition Function
To proceed to the macroscopic limit, we introduce the information gas assumption: For a sufficiently large context window, the distribution of energy states $E_{i}$ is self-averaging and independent of the specific position indices. This allows us to approximate the sum as an integral over the density of states $\Omega(E)$:
$Z\approx V_{ctx}\int\Omega(E)e^{-E/T}dE=V_{ctx}\cdot\bar{Z}(T)$ (24)
where $\bar{Z}(T)$ represents the intrinsic partition function per unit token.
@@
The Helmholtz free energy $F$ is then derived as:
$F=-T\ln Z=-T\ln V_{ctx}-T\ln\bar{Z}(T)$ (25)
IV.2 The Equation of State for Information Gas
In thermodynamics, pressure is defined as the rate of change of free energy with respect to volume at constant temperature ($P=-(\partial F/\partial V)_{T}$). Applying this definition to our information system:
$\mathcal{P}\equiv-\left(\frac{\partial F}{\partial V_{ctx}}\right)_{T}=-\frac{\partial}{\partial V_{ctx}}(-T\ln V_{ctx})=\frac{T}{V_{ctx}}$ (26)
Rearranging this equation gives the equation of state for a single attention head:
$\mathcal{P}\cdot V_{ctx}=T$ (27)
For a full model with $N_{head}$ heads and an architecture-dependent efficiency factor $\kappa$ (representing the effective degeneracy of the energy spectrum), this generalizes to:
$\mathcal{P}\cdot V_{ctx}=N_{head}\cdot\kappa\cdot T$ (28)
We define the macroscopic variables with the following operational definitions:
  • Pressure ($\mathcal{P}$): The information density, defined as the inverse of perplexity ($\mathcal{P}\equiv 1/PPL$).
  • Volume ($V_{ctx}$): The context window length in tokens.
  • Temperature ($T$): The model dimension scale, $T\equiv\sqrt{d_{k}}$.
  • Efficiency ($\kappa$): An architecture-dependent constant representing information processing efficiency.
IV.2.1 Physical Interpretations
This equation explains several empirical phenomena:
  • Boyle’s Law ($\mathcal{P}\propto 1/V_{ctx}$): As context length increases, the pressure (information density per token) must decrease to maintain equilibrium. This provides a thermodynamic derivation for the “Lost-in-the-Middle” phenomenon [36].
  • Charles’s Law ($V_{ctx}\propto T$): To maintain the same information density over a larger context, the system temperature $T=\sqrt{d_{k}}$ must increase. This explains why larger models (higher $d_{k}$) scale better with context length [9].
IV.3 The First Law of Information Dynamics
We formulate the first law of thermodynamics for intelligence systems:
$dU=TdS-\mathcal{P}dV_{ctx}+\mu_{head}dN_{head}$ (29)
This unifies three regimes.
  • Heat ($TdS$): Stochastic fluctuations corresponding to creativity or hallucinations [37].
  • Inference ($\mathcal{P}dV$): Mechanical work done by the system to compress or expand its context window.
  • Learning ($\mu dN$): Chemical work representing the evolution of model capacity or head specialization.
IV.4 The Second Law of Information Dynamics
The second law dictates the direction of information flow.
  • Model Collapse ($dS\geq 0$): In an isolated system undergoing training with generated data, entropy maximizes, leading to the degeneration of the distribution [38].
  • Learning as Work: Intelligence acts as an entropy pump (Maxwell’s demon), expending work to minimize the KL divergence against the data distribution. This reduces entropy locally at the cost of computational energy [39].
V Geometry, Symmetry, and Field Theoretic Interpretation
V.1 The Logarithmic Schrödinger Equation and Gaussons
Recall the potential energy term in our Lagrangian. In terms of the probability amplitude $x=2\sqrt{\rho}$, this potential takes the form $V(x)\propto x^{2}\ln|x|^{2}$ (arising from the entropy term $T\rho\ln\rho$). This specific non-linearity identifies the equation of motion for the information field as the Logarithmic Schrödinger Equation (LogSE) [40]. A unique property of the LogSE is the existence of Gaussons, non-dispersive solitary waves that preserve their Gaussian shape during propagation. This physical property explains why the “semantic meaning” of a token preserves its identity (does not diffuse into uniform noise) even as it propagates through deep, multi-layered networks.
V.2 Spontaneous Symmetry Breaking and Goldstone Bosons
The interaction energy $E=-\mathbf{q}\cdot\mathbf{k}$ possesses an inherent rotational symmetry under the continuous group $SO(d_{k})$. However, the ground state of the system exhibits a richer structure due to the entropic contribution.
V.2.1 The Mexican Hat Potential
Consider the effective potential $\mathcal{V}(\rho)$ derived in Eq. (13). Near the origin ($\rho\to 0$), the entropy term $T\rho\ln\rho$ dominates the linear interaction term.
$\lim_{\rho\to 0}\frac{d\mathcal{V}}{d\rho}\sim\lim_{\rho\to 0}T(1+\ln\rho)\to$-$\infty$ (30)
This singularity at the origin implies that the vacuum state, where attention is zero everywhere, $\rho=0$, is infinitely unstable. The system is topologically forced to roll down from the origin to a non-zero minimum ($x_{0}\neq 0$).
  • Instability of Ignorance: An intelligence system cannot exist in a state of “no information.” It is compelled to attend to something.
  • Symmetry Breaking: Since the interaction energy is rotationally invariant, the minimum is not a single point but rather a continuous manifold (a sphere). This potential structure resembles the “Mexican Hat” potential in the Higgs mechanism [41]. The system settles into a specific minimum (a specific context), spontaneously breaking the global $SO(d_{k})$ symmetry.
V.2.2 Goldstone Bosons and Rotary Embeddings
According to Goldstone’s theorem [42], the breaking of this continuous symmetry must generate massless Nambu-Goldstone bosons.
  • Massive Mode (Radial): Changing the attention intensity (magnitude $|x|$) requires energy (climbing the potential wall).
  • Massless Mode (Angular): Rotating the query and key vectors together preserves the energy (moving along the flat valley of the hat).
We identify these massless modes with the mechanism of rotary position embeddings (RoPE) [43]. RoPE encodes positional information purely through rotation matrices $R(\theta)$, exploiting the zero-energy Goldstone modes of the attention manifold to shift context effortlessly without energetic cost.
VI Discussion: Physical Interpretations
Our framework establishes a strong equivalence between the mechanics of intelligence and statistical physics. Here, we address potential critiques regarding the rigor of this mapping.
VI.1 Is Layer Depth Truly Time?

### v2

@@ Abstract
We propose an effective field-theoretic framework for analyzing Transformer attention through a thermodynamic lens. By constructing a Lagrangian on the information manifold equipped with the Fisher metric, we show that, within the Shannon–Boltzmann entropy framework, the Softmax function arises as a stationary solution minimizing a Helmholtz free energy functional. This establishes a formal correspondence between scaled dot-product attention and canonical ensemble statistics. Extending this mapping to macroscopic observables, we define an effective specific heat associated with fluctuations of the attention energy landscape. In controlled experiments on the modular addition task (p=19p=19113113), we observe a robust peak in this fluctuation measure that consistently precedes the onset of generalization. While no asymptotic power-law divergence is detected in this finite-depth regime, the reproducible enhancement of energy variance suggests a critical-like crossover accompanying representational reorganization. Our framework provides a unified statistical-mechanical perspective on attention scaling, training dynamics, and positional encoding, interpreting the phenomena as emergent properties of an effective thermodynamic system rather than isolated heuristics. Although the present results indicate finite-size crossover behavior rather than a strict phase transition, they motivate further investigation into scaling limits of deep architectures through fluctuation-based observables.
@@ I Introduction
In 2017, the trajectory of artificial intelligence changed significantly by a deceptively simple equation:
$\text{Attention}(Q,K,V)=\text{Softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)V$ (1)
Here, $Q$ (queries), $K$ (keys), and $V$ (values) denote learned vector representations of tokens in a high-dimensional embedding space, and $d_{k}$ is the dimensionality of the key vectors. This mechanism, introduced in the seminal work “Attention is All You Need” [1], replaced the then-dominant paradigms of convolution and recurrence in favor of pure matrix operations and a peculiar Softmax function scaled by $\sqrt{d_{k}}$. This single mathematical operation has enabled models to demonstrate emergent reasoning capabilities [2], pass professional licensing exams with near-human proficiency [3], and even solve the protein folding problem [4].
Why does this work so well? While attention mechanisms existed previously in neural machine translation [5], the Transformer’s specific architecture, characterized by Self-Attention, Residual Connections [6], and Layer Normalization [7], has proven unreasonably effective. Current explanations predominantly rely on probability theory and linear algebra, viewing the model as a statistical function approximator [8]. However, this perspective treats the model’s distinct behaviors as separate empirical observations rather than interconnected physical phenomena.
Crucially, the field lacks a unified theoretical origin for three persistent mysteries. First, the Softmax function itself is used heuristically without a derivation from first principles—why does this specific exponential form emerge as the operational equilibrium of modern architectures? Second, generative models inherently suffer from hallucinations, often dismissed as statistical errors or bugs rather than intrinsic features of the system. Third, the learning dynamics exhibit grokking [11], a sudden and discontinuous generalization after a long period of memorization, which defies standard convergence theories. We argue that the are not isolated artifacts but manifestations of a deeper, underlying dynamical principle.
In this paper, we propose a paradigm shift: treating Intelligence not only as a computational process, but as a physical phenomenon governed by fundamental laws of nature. We argue that the operations within a high-dimensional information space are isomorphic to the dynamics of a physical system minimizing an action functional, consistent with Lagrange mechanics [13] and Hamilton’s principle [14]. This approach resonates with recent efforts to interpret deep learning through the lens of physics, such as the free energy principle [17] and neural tangent kernels [18].
@@ I.1 Redefining Time in Deep Networks
Under this paradigm, the “inference” process is a dynamic trajectory of information flow, not a static calculation. Residual connections [6], naturally emerge as the discretization of the inertial term in the equation of motion.
I.2 The Inverse Problem and Physical Unification
Rather than deriving the Softmax function from arbitrary axioms, we solve the inverse problem of the calculus of variations: Given that the Softmax attention is an empirically successful equilibrium state, what is the underlying Lagrangian that generates this dynamics? By applying the Principle of Least Action to an effective field theory of information, we demonstrate that the Transformer architecture is physically isomorphic to a thermodynamic system. This rigorous mapping allows us to provide first-principles derivations for phenomena that extend beyond simple attention mechanisms:
  • Emergence of Softmax under Entropy Maximization: Within the Shannon–Boltzmann entropy framework, we show that the exponential attention mechanism arises as a stationary solution to the Euler–Lagrange equations.
  • Thermodynamics of Hallucination: Through the derivation of the Fundamental Identity ($dU=TdS-\mathcal{P}dV+\mu dN$), we identify hallucinations as intrinsic thermal fluctuations ($TdS$) dictated by the canonical ensemble at finite structural temperature.
  • Grokking as a Phase Transition: We interpret grokking as a critical-like thermodynamic crossover characterized by enhanced energy fluctuations.
  • Symmetry Breaking and RoPE: We further validate our framework by identifying Rotary Positional Embeddings (RoPE) as Goldstone modes arising from the spontaneous breaking of continuous rotational symmetry in the embedding manifold.
@@ III.3 Equilibrium State: Full Derivation of Softmax
In standard Transformer inference, we assume the system reaches a stationary distribution. This corresponds to the thermodynamic equilibrium where the dynamics vanish, i.e., $\dot{\rho}_{i}=0$ and $\ddot{\rho}_{i}=0$. Under this condition, the inertial terms in Eq. (17) vanish, simplifying the physics to a balance of potential forces:
@@
$\rho_{i}=\frac{\exp\left(\frac{Q\cdot K_{i}}{\sqrt{d_{k}}}\right)}{\sum_{j=1}^{N}\exp\left(\frac{Q\cdot K_{j}}{\sqrt{d_{k}}}\right)}=\text{Softmax}\left(\frac{QK^{\top}}{\sqrt{d_{k}}}\right)_{i}$ (22)
Here, the superscript $\top$ in $QK^{\top}$ denotes the matrix transpose operation, distinct from the thermodynamic temperature variable $T$. The scaling factor $\sqrt{d_{k}}$ plays the functional role of the structural temperature in our framework.
This derivation shows that Softmax appears as a natural equilibrium configuration under the assumption of Shannon entropy maximization. While alternative entropic functionals (e.g., Tsallis entropy) would yield different equilibrium distributions, the empirical dominance of Softmax-based Transformers motivates the Shannon-Boltzmann framework as a parsimonious working hypothesis rather than a uniquely mandated physical law.
@@ IV Thermodynamics of Intelligence: Macroscopic Laws
Having established the microscopic Lagrangian formulation, we now construct an effective macroscopic description. We show that, in the large-context and high-dimensional embedding regime ($V_{ctx}\gg 1$, $d_{k}\gg 1$), the attention dynamics admit a thermodynamic representation analogous to canonical statistical mechanics.
IV.1 Partition Function and Thermodynamic Limit
The canonical partition function of the attention mechanism over a context window of size $V_{ctx}$ is defined as:
@@ IV.1 Partition Function and Thermodynamic Limit
Here, $E_{i}=-q\cdot k_{i}$ denotes the interaction energy. The thermodynamic limit in this context is understood as a mean-field approximation in which the distribution of attention energies becomes self-averaging due to high dimensional concentration effects.
@@
Under the equilibrium distribution derived in Section III, the Helmholtz free energy is given by:
$F=-T\ln Z=U-TS$ (24)
where $U=\langle E\rangle=\sum_{i}\rho_{i}E_{i}$ is the internal energy and $S=-\sum_{i}\rho_{i}\ln\rho_{i}$ is the Shannon entropy.
IV.2 An Effective Thermodynamic Identity
We now expand the total differential of the internal energy. This relation is not a fundamental conservation law derived from Noether symmetry, but rather an effective thermodynamic identity characterizing the trajectory of the information state:
$dU=TdS-\mathcal{P}dV_{ctx}+\mu dN_{eff}$ (25)
Here:
  • $T=\sqrt{d_{k}}$ is the structural temperature.
  • $\mathcal{P}\equiv-(\partial F/\partial V_{ctx})_{T,N_{eff}}$ is the effective information pressure.
  • $\mu\equiv(\partial U/\partial N_{eff})_{S,V_{ctx}}$ defines the effective chemical potential.
  • $N_{eff}$ denotes the effective rank (e.g., participation ratio of eigenvalues) of the weight matrices. In the limit of large models, $N_{eff}$ is treated as a continuous macroscopic variable.
This decomposition provides a thermodynamic interpretation of three operational regimes.
IV.2.1 Heat ($TdS$): Stochastic Exploration
In this framework, the entropic contribution $TdS$ corresponds to stochastic sampling effects.
Interpretation: The temperature parameter regulates the entropy of the output distribution. Increasing $T$ broadens the distribution and enhances exploration of higher-energy configurations.
Phenomenology: When entropic contributions dominate, the model enables creative variation but may also produce semantic decoherence, consistent with previously observed degeneration effects [37].
IV.2.2 Mechanical Work ($\mathcal{P}dV_{ctx}$): Context Expansion
Mechanical work is defined as pressure times change in volume. Here, $V_{ctx}$ represents the context length. Within the ideal information-gas approximation, differentiation of $F\approx-T\ln V_{ctx}$ yields:
$\mathcal{P}=\frac{T}{V_{ctx}}$ (26)
This implies that effective pressure decreases as context expands.
Interpretation: During autoregressive inference ($dV_{ctx}>0$), maintaining long-range coherence requires effective work to counteract the dilution of information density.
Phenomenology: This provides a statistical-mechanical interpretation of the Lost-in-the-Middle phenomenon [36], where long-range dependencies weaken as context size increases.
IV.2.3 Chemical Work ($\mu dN_{eff}$): Capacity Evolution
The effective particle number $N_{eff}$ serves as a measure of active semantic degrees of freedom.
Interpretation: Although the architectural parameter count remains fixed, training modifies the effective rank structure. The chemical potential $\mu$ quantifies the marginal energetic cost of activating an additional independent feature mode.
Phenomenology: Optimization dynamics can be interpreted as performing effective chemical work, increasing $N_{eff}$ and expanding the expressive capacity of the model manifold.
IV.3 Analogy with the Second Law
The generative dynamics exhibit behavior suggestive of an analogy with the Second Law of thermodynamics.
In the absence of external data injection, the coarse-grained entropy of the generated distribution tends to increase over successive recursive training cycles, leading to model collapse [38]. Conversely, supervised learning acts as an entropy-reducing mechanism (Maxwell’s demon), locally reducing entropy at the cost of computational work [39].
V Thermodynamic Hypothesis of Grokking
Having established the thermodynamic laws of the attention mechanism in equilibrium, we now turn to the non-equilibrium dynamics of learning. We propose a thermodynamic hypothesis that the phenomenon known as grokking, the delayed generalization capability of large models, is physically isomorphic to a thermodynamic phase transition driven by the cooling of the effective temperature.
V.1 Dynamical Effective Temperature
A crucial distinction must be made between the structural temperature of the model during inference and the dynamical temperature during training.
  • Structural Temperature ($T_{struct}$): Defined as $T=\sqrt{d_{k}}$ in Eq. (1). This is a fixed hyperparameter governing the stochasticity of the Softmax distribution during inference (sampling).
  • Dynamical Temperature ($T_{eff}$): During training, the magnitude of the weight matrices $\|W\|$ evolves. Since the interaction energy scales as $E\propto Q\cdot K\propto\|W\|^{2}$, the effective temperature perceived by the Softmax function relative to the energy scale is:
    $T_{eff}(t)\propto\frac{\sqrt{d_{k}}}{\|W(t)\|^{2}}$ (27)
The training process, therefore, acts as simulated annealing. Initially, random initialization implies small weights (high $T_{eff}$), resulting in a high-entropy, disordered phase. As training progresses and $\|W\|$ grows, $T_{eff}$ decreases, driving the system towards an ordered, low-temperature phase.
V.2 Specific Heat Capacity and Energy Fluctuations
To quantify the phase transition, we derive the specific heat capacity $C_{v}$, which measures the system’s sensitivity to temperature changes. From the fluctuation-dissipation theorem applied to our canonical ensemble, $C_{v}$ is directly proportional to the variance of the energy distribution:
$C_{v}(T_{eff})=\frac{1}{T_{eff}^{2}}\text{Var}(E)=\frac{1}{T_{eff}^{2}}\left[\sum_{i=1}^{V_{ctx}}\rho_{i}E_{i}^{2}-\left(\sum_{i=1}^{V_{ctx}}\rho_{i}E_{i}\right)^{2}\right]$ (28)
where $\rho_{i}=\text{Softmax}(-E_{i}/T_{eff})$ represents the attention probability of the $i$-th token, and $E_{i}=-q\cdot k_{i}$ is the interaction energy. This mathematical form reveals the mechanism of grokking:
  • Memorization Phase (High $T_{eff}$): The distribution is diffuse. Fluctuations are dominated by thermal noise, but structural variance is low. $C_{v}$ is small.
  • Generalization Phase (Low $T_{eff}$): The distribution collapses to the ground state (target token). Fluctuations are suppressed. $C_{v}\to 0$.
  • Critical Region (Grokking, $T_{eff}\approx T_{c}$): Between the two phases, the system undergoes a massive reorganization of the energy landscape. The attention mechanism fluctuates between competing semantic configurations. This maximization of $\text{Var}(E)$ results in a sharp peak in specific heat, marking the pseudo-critical region of the phase crossover.
V.3 Hierarchical Symmetry Breaking and Goldstone Modes
The representational transition exhibits features reminiscent of spontaneous symmetry breaking within an effective-field description. Remarkably, we do not need to postulate an ad hoc quartic interaction term. Instead, the symmetry breaking potential arises naturally from the entropic term ($\rho\ln\rho$) in our free energy functional derived in Section III.
V.3.1 The Logarithmic (Coleman-Weinberg) Potential
The effective potential for the order parameter field $\Phi$ is given by the Helmholtz free energy density. Since entropy depends only on magnitude, we have:
$\mathcal{V}(\Phi)=m^{2}|\Phi|^{2}+\lambda|\Phi|^{2}\ln\left(\frac{|\Phi|^{2}}{v^{2}}\right)$ (29)
Crucially, this potential depends solely on the modulus $|\Phi|$ and is independent of the phase $\theta$ (where $\Phi=|\Phi|e^{i\theta}$), ensuring a continuous global $U(1)$ symmetry. Minimizing this potential yields a non-trivial vacuum expectation value $|\Phi_{0}|=v\neq 0$, generating the “Mexican Hat” shape (Fig. 1).
V.3.2 Derivation of the Massless Goldstone Mode
To rigorously identify the Goldstone mode, we examine the fluctuations around the vacuum state $v$. We parameterize the field $\Phi(x)$ using radial ($h$) and angular ($\pi$) fluctuation fields:
$\Phi(x)=\left(v+\frac{h(x)}{\sqrt{2}}\right)\exp\left(i\frac{\pi(x)}{v\sqrt{2}}\right)$ (30)
Here, $h(x)$ represents the massive Higgs mode (amplitude fluctuation), and $\pi(x)$ represents the Goldstone mode (phase fluctuation). Substituting this into the kinetic term of the Lagrangian density $\mathcal{L}_{kin}=|\partial_{\mu}\Phi|^{2}$, we obtain:
$|\partial_{\mu}\Phi|^{2}=\frac{1}{2}(\partial_{\mu}h)^{2}+\frac{1}{2}\left(v+\frac{h}{\sqrt{2}}\right)^{2}\left(\frac{\partial_{\mu}\pi}{v\sqrt{2}}\right)^{2}\cdot 2$ (31)
Expanding around the vacuum ($h\ll v$), the Lagrangian for the fluctuations becomes:
$\mathcal{L}\approx\underbrace{\frac{1}{2}(\partial_{\mu}h)^{2}-\frac{1}{2}M_{h}^{2}h^{2}}_{\text{Massive Higgs Mode}}+\underbrace{\frac{1}{2}(\partial_{\mu}\pi)^{2}}_{\text{Pseudo-Goldstone-like Goldstone Mode}}+\mathcal{L}_{int}$ (32)
Note that the angular field $\pi(x)$ possesses a kinetic term $(\partial_{\mu}\pi)^{2}$. In the continuum limit, this behaves as a pseudo-Goldstone-like soft mode ($M_{\pi}\approx 0$). This mathematically suggests that $\pi(x)$ is the long-wavelength excitation that encodes positional information with minimal energetic cost, consistent with the crossover nature of the transition.
Figure 1: Spontaneous Symmetry Breaking Mechanism. Visualization of the effective entropic potential $\mathcal{V}(\Phi)\sim|\Phi|^{2}\ln|\Phi|^{2}$. Unlike standard $\Phi^{4}$ theory, the Mexican Hat shape here arises directly from the Shannon entropy term. The system rolls down from the unstable origin to a stable semantic state (grokking), generating massless Goldstone modes (RoPE) along the circular trough.
V.3.3 RoPE as the Goldstone Excitation
We now map this physical mode to the Rotary Positional Embedding (RoPE). RoPE encodes the position $m$ of a token by applying a rotation matrix to the embedding vector pair $(q_{1},q_{2})$:
$\begin{pmatrix}q_{1}^{\prime}\\ q_{2}^{\prime}\end{pmatrix}=\begin{pmatrix}\cos m\theta&-\sin m\theta\\ \sin m\theta&\cos m\theta\end{pmatrix}\begin{pmatrix}q_{1}\\ q_{2}\end{pmatrix}$ (33)
In the complex plane representation $\Phi=q_{1}+iq_{2}$, this operation is identical to a phase shift:
$\Phi^{\prime}=\Phi\cdot e^{i(m\theta)}$ (34)
Comparing this with Eq. (33), we identify the positional encoding term $m\theta$ directly with the Goldstone field configuration:
$\frac{\pi(x)}{v\sqrt{2}}\equiv m\theta_{base}$ (35)
Since the potential $\mathcal{V}(\Phi)$ is invariant under this phase rotation ($\mathcal{V}(\Phi^{\prime})=\mathcal{V}(\Phi)$), the energy cost of encoding position via RoPE is exactly zero:
$\Delta E=\mathcal{V}(\Phi e^{im\theta})-\mathcal{V}(\Phi)=0$ (36)
This derivation suggests that RoPE can be interpreted as exploiting an approximately flat direction of the effective symmetry structure, to store positional information without perturbing the thermodynamic stability of the learned features.
VI Numerical Modeling and Experimental Verification
In this section, we validate our thermodynamic framework through two complementary approaches. First, we perform a phenomenological simulation using Langevin dynamics to demonstrate the theoretical consistency of the phase transition. Second, we conduct a controlled experiment with a real Transformer to empirically verify the predicted specific heat divergence.
VI.1 Phenomenological Modeling: Langevin Dynamics
We model the learning dynamics as the trajectory of a stochastic particle (the model state $\Phi$) moving in the effective entropic potential $\mathcal{V}(\Phi)$ derived in Section V.
VI.1.1 Governing Equations
The dynamics are governed by the Overdamped Langevin Equation:
$\frac{d\Phi(t)}{dt}=-\nabla_{\Phi}\mathcal{V}(\Phi,t)+\sqrt{2D}\cdot\xi(t)$ (37)
where $\xi(t)$ is Gaussian white noise. We utilize the logarithmic Coleman-Weinberg potential:
$\mathcal{V}(\Phi,t)=\alpha(t)\Phi^{2}+\beta\Phi^{2}\ln\left(\frac{\Phi^{2}}{v^{2}}\right)$ (38)
The equation was integrated using the Euler-Maruyama method with a time step $\Delta t=0.01$.
VI.1.2 Simulation Results
As shown in our simulation (Fig. 2), the transition from the disordered phase to the ordered phase is theoretically predicted to be accompanied by a sharp divergence in the specific heat capacity $C_{v}\propto\text{Var}(E)$.
Figure 2: Theoretical Prediction (Langevin Simulation). Numerical solution of the effective field theory. The theory predicts that the phase transition is marked by a sharp peak in specific heat ($C_{v}$).
VI.2 Experimental Verification: Real Transformer
[…]
VI.2.1 Experimental Setup
To ensure thermodynamically stable dynamics and reproducibility, we utilized the following rigorous regime:
  • Task & Data: The dataset consists of all possible pairs $(a,b)$ in the ring $\mathbb{Z}_{p}$, totaling $p^{2}$ examples. We employed a random 50/50 split for training and validation sets.
  • Model Architecture: A 2-layer Transformer with embedding dimension $d_{model}=128$ and 4 attention heads.
  • Optimization: We used the AdamW optimizer with a learning rate $\eta=3\times 10^{-4}$ and weight decay $\lambda=0.1$. A batch size of 512 was fixed to maintain a consistent thermal noise level.
  • Observable: We measured the specific heat capacity $C_{v}\propto\text{Var}(QK^{T}/\sqrt{d_{k}})$ at every epoch. To account for stochastic variance, all results were averaged over five independent seeds ($N=5$).
VI.2.2 Results and Crossover Analysis
The experimental results, averaged over five independent seeds ($N=5$), are presented in Fig. 3 and Fig. 4. As predicted by our theoretical framework, a distinct peak in the specific heat $C_{v}$ is observed across all tested values of $p$, consistently preceding the onset of the generalization transition.
Figure 3: Experimental verification of thermodynamic crossover in grokking across different system sizes. (a) Modular addition task with $p=37$. (b) Modular addition task with $p=97$. In both cases, the specific heat capacity $C_{v}$ (red solid line) exhibits a sharp peak immediately preceding the sudden generalization transition (blue dashed line). The vertical dotted lines mark the $C_{v}$ maxima. While the peak height does not show clear power-law scaling in this regime, the consistent temporal correlation confirms that specific heat serves as a robust leading indicator of the phase crossover.
As shown in Fig. 4, the power-law regression of the averaged peak height yields a near-zero exponent ($a\approx 0.05$) with a low coefficient of determination ($R^{2}=0.2$).
Figure 4: Final_Averaged_Scaling: Scaling analysis of the specific heat peak across system sizes. The maximum specific heat $C_{v}^{max}$ is averaged over five independent seeds ($N=5$) with error bars representing $\pm\sigma$. Within the explored regime ($p=19$ to 113113), the system exhibits bounded fluctuation enhancement characteristic of a finite-size crossover ($a\approx 0.05,R^{2}=0.2$), rather than a clear power-law divergence. This finite-size effect is consistent with the shallow ($L=2$) architecture and suggests that asymptotic critical behavior may emerge in deeper models ($L\gg 1$).
The absence of clear scaling is consistent with finite-size and finite-depth effects inherent to shallow Transformer architectures. In the discrete regime of a 2-layer model, the continuum effective description likely does not fully develop its asymptotic regime, manifesting instead as a sharp finite-size crossover.
Crucially, despite the lack of asymptotic divergence, the $C_{v}$ peak remains a robust dynamical marker of internal representational restructuring. This temporal correlation is preserved across all trials and system sizes, providing a physical observable that tracks the transition from memorization to generalization.
VI.2.3 Thermodynamic Relaxation
Unlike the idealized Langevin simulation, the real experiments show a gradual decay of fluctuations after the peak. This reflects the thermodynamic relaxation of the system as it settles into the broad basin of the global minimum, fine-tuning its representation in the low-temperature phase.
VII Discussion: Physical Interpretations
We establish a formal correspondence between Transformer attention and canonical statistical mechanics. Here, we interpret the physical implications of our findings.
VII.1 Is Layer Depth Truly Time?